The supply function for a good can be written as Q = 2P + 10, where Q is the quantity supplied in Kilos and P is the Price in dollars.

The price falls from $15 to $10.

Calculate the price elasticity of supply (PES).




1
Expert's answer
2021-11-30T17:42:09-0500

"P_1 = 15 \\
Q_1 = 2 \times 15 + 10 = 40 \\
P_2 = 10 \\
Q_2 = 2 \times 10 + 10 = 30 \\
PES = \\frac{\%ΔQ}{\%ΔP} \\
\%ΔQ = \frac{(40-30) \times 100 \; \%}{40} = 25 \; \% \\
\%ΔP = \frac{(10-15) \times 100 \; \%}{15} = -33.33 \; \% \\
PES = \frac{25\; \%}{-33.33\; \%} = -0.75"